# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

### v4

@@ II.1 The Hamiltonian of the quantum-classical system
II.2 Mixed quantum-classical equilibrium density
The equilibrium density corresponding to the Hamiltonian of Eq. (1) can be written as
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}},$ (9)
where $\beta=(\text{k}_{\text{B}}T)^{-1}$, $\text{k}_{\text{B}}$ is the Boltzmann constant, and the dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables ${\bf{Q}}$ is here explicitly highlighted. This result is nothing but the real-space representation of that obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis. The result of Eq. (9) can be obtained by starting from an Hamiltonian where both sets of particles are considered to be quantum. Then, one performs a Wigner transform over the $N_{{\text{mm}}}$ particles only and eventually a semi-classical limit is obtained by letting the ratio of masses $(m/M)^{1/2}$ tends towards 0. Conserving only the zeroth order term of this semi-classical limit yields to Eq. (9). This expansion in terms of the ratio of the masses is natural since the quantum particles, typically electrons, are usually much lighter than the classical ones, typically atoms or molecules. The original derivation of Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) used the Liouvillian imaginary time evolution equation of the density matrix together with the Poisson-brakets operatorsİmre et al. (1967), which is a well-known formalism in the quantum dynamics communityKapral and Ciccotti (1999); Nielsen, Kapral, and Ciccotti (2001); Kapral (2006, 2015). However, we provide in the supplementary materials an alternative derivation which involves a similarity transformation and resorts to the Baker-Campbell-Hausdorf expansion to express the Wigner transform Hamiltonian as a series of nested commutators, which might appear more familiar to the ”static” quantum chemistry community.
In Eq. (9), the normalization is the partition function of the QM/MM system which is expressed as
$Z^{\text{tot}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}}\int\text{d}{\bf{Q}}\text{d}{\bf P}\!\!\int\!\text{d}{\bf{q}}$ ${\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}$ (10)
$e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}.$
The corresponding Helmholtz free energy is given by
$F_{0}^{\text{tot}}=-\text{k}_{\text{B}}T\log(Z^{\text{tot}}).$ (11)
While the expressions of Eqs. (9) and (10) are given in real-space, they can easily be written in the adiabatic basis, where one explicitly sees that the density matrix is diagonal as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001).
@@ II.2 Mixed quantum-classical equilibrium density
Although the expression of the equilibrium density in Eq. (9) is relatively simple, it calls some comments. Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically. Therefore it is worth splitting the equilibrium density in terms of ${\bf{Q}}$ and ${\bf P}$ dependence
@@
$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg{(}\frac{2\pi M}{\beta h^{2}}\bigg{)}^{\frac{3}{2}N_{{\text{mm}}}},$ (15)
@@
$F_{0}=-\text{k}_{\text{B}}T\log(Z)$ (20)
is the interacting part of the Helmholtz free energy of the QM/MM system.
Secondly, the object $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ in Eq.(13) can be expressed as the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
@@
It is noteworthy that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ defined in Eq. (22),(23) is a positive definite operator and that $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$ of Eq. (25) is also positive definite and normalized to unity. Therefore, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are simply a special case of more general density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow us in Sec. II.3 to propose a variational formulation of the Helmholtz free energy for a QM/MM system.
Additionally, this decomposition of the equilibrium density as a product of a purely quantum equilibrium density and a purely classical equilibrium density could lead to the somehow strange conclusion that there is no correlation between the quantum and classical systems, even if the two systems interact. Nevertheless, one must pay attention to the fact that the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. Therefore, the two systems are indeed correlated.
Thirdly, an obvious illustration of the meaning of the equilibrium density of Eq. (13) is that corresponding to the Born-Oppenheimer approximation. In the latter case, the light particles described by ${\bf{q}}$ are electrons and the heavy particles described by ${\bf{Q}}$ are the classical nuclei. There, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, there is no external potential both for the QM and MM parts, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is the nuclei-electron attraction, and ${\bf W}_{{\text{mm}}}({\bf{Q}})$ is the Coulomb repulsion between the nuclei, such that the operator happening in the exponential of Eq. (13) is the Born-Oppenheimer Hamiltonian.
A follow up on this comment is that, as mentioned earlier, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is diagonal in the adiabatic basis, which is equivalent to say that non-adiabatic couplings are completely neglected. In the context of ab initio MD, this implies that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ can be obtained by averaging, over all possible adiabatic electronic states, the classical trajectories of nuclei computed from the forces of a given adiabatic electronic state. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature concerned with going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM treatments (see for instance Ref. Kapral, 2015 for a review), but the treatment of non adiabatic couplings goes beyond our QM/MM scheme.
II.3 Reformulation of the Helmholtz free energy for a QM/MM system within a variational framework
@@ II.3.1 Variational formulation
As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we want to reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq. (21) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq. (20). Moreover, since the ultimate goal is to formulate a variational principle involving both quantum and classical densities, it is natural to use the decomposition of Eq. (21).
@@
where the operator $\hat{{\boldsymbol{\rho}}}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
@@
and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable. Therefore, ${\boldsymbol{\rho}}$ and $\boldsymbol{p}$ must be normalized to be density matrices and probability measures, respectively, i.e
@@
$\text{Tr}\{\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\}=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (31)
@@
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (32)
where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has was accounted for in $F_{{\bf P}}$)
@@
$-\text{k}_{\text{B}}T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}},\boldsymbol{p}),$ (35)
@@
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=F_{0}=-\text{k}_{\text{B}}T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (36)

### v5

@@ II.1 The Hamiltonian of the quantum-classical system
II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
We start by defining the exact equilibrium density matrix (EDM) of the QM/MM system in the canonical ensemble. This quantity is the analogue of the equilibrium density matrix of a purely quantum system or, equivalently, of the equilibrium probability density of a purely classical (MM) system.
With the Hamiltonian of Eq. (1), the EDM takes the following form
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ $=\frac{{\langle{\bf{q}}|}e^{-\beta\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})}{|{\bf{q}}^{\prime}\rangle}}{Z^{\text{tot}}}$ (9)
$=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}}.$
In Eq. (9) $\beta=T^{-1}$ in atomic units, and $|{\bf{q}}\rangle$} and $|{\bf{q}}^{\prime}\rangle$} are position eigenstates for the $N_{{\text{qm}}}$ quantum variables. The dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables, $\bf{Q}$}, has been made explicit. Because Eq. (9) involves the exponential of a quantum operator, the QM/MM EDM is itself a quantum operator, namely a quantum density matrix (i.e. positive definite and of trace unity). However, it is parametrized by the classical variables $({\bf{Q}},{\bf P})$ and therefore the QM system depends on the configuration of the classical particles. To understand the physical meaning of the EDM, it is instructive to analyse its diagonal elements, i.e. when ${\bf{q}}={\bf{q}}^{\prime}$. The quantity $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}})$ gives the probability of finding the quantum particles at positions ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$ while the classical particles have positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momenta ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$. Off-diagonal elements encode purely quantum effects, such as nonlocality, and are necessary to compute the action of quantum operators—for example, the kinetic energy operator, which involves a Laplacian in the quantum subsystem.
A pictorial illustration of the physical meaning of the equilibrium density of Eq. (9) is provided by the Born-Oppenheimer description of a molecule, i.e. QM electrons and classical nuclei. In that case, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, with no external potential for either the QM or MM parts. The operator $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ represents the nuclei-electron attraction, while ${\bf W}_{{\text{mm}}}({\bf{Q}})$ represents the Coulomb repulsion between the nuclei. Thus, $\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})$ entering in the exponential of Eq. (9) corresponds the Born-Oppenheimer Hamiltonian.
From a more formal point of view, it is worth noticing that Eq. (9) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis, i.e. in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. In this basis, the density matrix is diagonal , as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001). In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian. Therefore, Eq. (9) ignores the non-adiabatic couplings between these states. In the context of ab initio MD, this implies that $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ can be obtained from classical nuclear trajectories, with forces averaged over all adiabatic electronic states weighted by the corresponding Boltzmann factors. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature addressing methods going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM systems (see for instance Ref. Kapral, 2015 for a review), but treatment of non adiabatic couplings is beyond the scope of this paper.
[…]
In Eq. (9), the normalization corresponds to the partition function of the QM/MM system which is expressed as
$Z^{\text{tot}}=\frac{1}{N_{{\text{mm}}}!}$ $\int\text{d}{\bf{Q}}\text{d}{\bf P}\,e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}$ (10)
$\int\!\text{d}{\bf{q}}{\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}.$
The corresponding Helmholtz free energy is therefore given by
$F_{0}^{\text{tot}}=-T\log(Z^{\text{tot}}).$ (11)
Any equilibrium property $\langle O\rangle_{\text{eq}}$ can be obtained from derivatives of $Z^{\text{tot}}$. Alternatively, $\langle O\rangle_{\text{eq}}$ can be computed as the trace of the product of $\mathcal{M}$ and the operator $\hat{O}$, which may depend on both quantum and classical variables
$\langle O\rangle_{\text{eq}}$ $=\int\text{d}{\bf{Q}}\text{d}{\bf P}\,\int\!\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}\hat{O}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}}^{\prime},{\bf{q}})$ (12)
$\equiv\text{Tr}\{\hat{O}\hat{\mathcal{M}}\}$.
@@ II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
Although the expression of the equilibrium density in Eq. (9) is relatively simple, it benefits from further comments. Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically. Therefore, it is convenient to separate the equilibrium density according to its dependence on $\bf{Q}$} and $\bf P$}
@@
$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg(\frac{2\pi M}{\beta}\bigg)^{\frac{3}{2}N_{{\text{mm}}}},$ (16)
@@
$F_{0}=-T\log(Z).$ (21)
$F_{0}$ represents the contribution to the Helmholtz free energy of the QM/MM system arising from particle interactions. Computing $F_{0}$ from Eq. (21) is infeasible for realistic systems, as it requires evaluating Eq. (17). The aim of this paper is to develop an alternative, more practical, route to $F_{0}$.
The EDM $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ of Eq.(14) can be decomposed into the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
@@
It is noteworthy that, for each value of $\bf{Q}$}, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})$ defined in Eq. (23) and (24) is a density matrix, i.e. positive definite and of unity norm. Similarly, $\boldsymbol{p}_{\text{eq}}$ of Eq. (26) is a probability distribution. Therefore, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are special cases of more general objects: density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec. II.3.
Last but not least, the decomposition of the equilibrium density of Eq. (22) into a product of a purely quantum equilibrium density and a purely classical equilibrium density might suggest that the quantum and classical subsystems are uncorrelated, even though they interact. Nevertheless, the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. The classical part, $\boldsymbol{p}_{\text{eq}}$, is in turn affected by the quantum degrees of freedom through its normalization, $Z^{{\text{qm}}}({\bf{Q}})$. To see this more explicitly, one can notice that the form of Eq. (26) can be rewritten as
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{1}{Z}e^{-\beta\big({{\bf{V}}}_{\text{mm}}({\bf{Q}})+U_{\text{eff}}({\bf{Q}})\big)},}$ (28)
where now the classical particles experience an additional the effective potential $U_{\text{eff}}({\bf{Q}})$ due to the presence of the QM system, which is defined as
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{\text{eff}}({\bf{Q}})=-\frac{1}{\beta}\log\big(Z^{{\text{qm}}}({\bf{Q}})\big).$} (29)
Therefore, the two systems are correlated.
II.3 Variational formulation of the Helmholtz free energy for a QM/MM system
Both classical and quantum DFT can be formulated through the Levy-Lieb constrained-search formalism. The first step is to express the quantum ground state energy and the classical Helmholtz free energy as a general functional minimization problem, defined over the many-body functions, namely the electronic density matrix and the probability distribution, respectively.
Since the goal of this work is to obtain a DFT formulation of the QM/MM problem, the first step is to reformulate the Helmholtz free energy of Eq. (21) as a functional minimization problem over QM/MM density matrices whose form generalize the EDM of Eq. (22). This is the aim of this section. To the best of our knowledge, this was never formalized before.
@@ II.3.1 Variational formulation
As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we can now reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq. (22) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq. (21).
@@
where the operator $\hat{{\boldsymbol{\rho}}\,}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also parametrically depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
@@
and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable. Therefore, $\hat{{\boldsymbol{\rho}}\,}$ and $\boldsymbol{p}$ must be normalized to unity in order to be acceptable density matrices and probability measures, respectively, i.e
@@
$\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (34)
@@
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (35)
where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has already been accounted for in $F_{{\bf P}}$)
@@
$-T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}),$ (38)
@@
$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=F_{0}=-T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}\,}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (39)
This variational formulation, is the analogue of the Gibbs or Von Neumann variational principles, but applied to a QM/MM system. As with any variational formulation, it naturally enables the construction of approximations by imposing constraints on the variables of minimization, the archetypal example being the mean-field approach, which is briefly discussed in Sec. III.3.